Find the thermo-emf developed in a copper-silver thermocouple when the junctions are kept at 0°C and 40°C. Use the data in table (33.1).

Grade:10

1 Answers

Jitender Pal
9 years ago
Sol. E = a base ABθ + b base ABθ^2 a base CuAg = a base CuPb – b base AgPb = 2.76 – 2.5 = 0.26 μv/°C b base CuAg = b base CuPb – b base AgPb = 0.012 – 0.012 μvc = 0 E = a base ABθ = (0.26 × 40) μV = 1.04 × 10^–5 V